The following balance sheet proportions belong to Shelbee Company. Shelbee can issue long-term bonds at current...

  1. The following balance sheet proportions belong to Shelbee Company. Shelbee can issue long-term bonds at current yield to maturity of 10%. It has just paid annual dividend of $2.00 per share and its dividends grow at a constant rate of 5%. Its shares trade at $25.00 per share. Shelbee’s marginal income tax rate is 20%.

Book Value %

Market Value %

Target Capital Structure %

Assets

100

100

100

Liabilities

40

30

25

Equity

60

70

75

  1. Calculate WACC for Shelbee.
  2. Assume Shelbee can raise capital up to $800,000 at the costs calculated above (in a). If more capital is needed, it has to issue equity and face floatation costs of 20%. What would be the WACC with new equity issue?
  3. Assume Shelbee has the following independent projects to evaluate.

Project

Cost $

IRR %

Green

300,000

14

Red

200,000

13

Orange

400,000

15

Purple

250,000

12

Which projects should it undertake and what is its capital budget?
Answer #1
a Calculation of WACC
For calculating target value rates are considered.
Ratio Cost WACC
Equity 0.75 13.40% 10.05
Debt 0.25 8% 2
12.05
Kd = Post-tax yield to maturity
      = (10*0.80) 8
Re = (D1/ Price)+g
0.134    or 13.4%
D1 =      2*1.05 2.1
Price 25
b New Price = 25 *0.80 20
(After floatation cost)
New Re 0.155 15.50%
Ratio Cost WACC
Equity 0.75 15.5 11.625
Debt 0.25 8 2
13.625
c IRR means an internal rate of return from the project that the company
will earn. Therefore the project with higher IRR should be undertaken provided
that it should exceed WACC
Project Cost IRR Preference
Green 300000 14 2
Red 200000 13 3
Orange 400000 15 1
Purple 250000 12 4
If the company has $800000 then it should invest in project orange and green.
The further capital raised should not be invested in project Red because IRR is 13
which is less than its cost i.e WACC
Therefore raising of further capital beyond $ 800000should not be recommended